I really need help on solving this word problem. It would be GREATLY appreciated if someone could help me!
- During one month, a homeowner used 500 units of electricity and 100 units of gas for a total utility bill of $88. The next month, 400 units of electricity and 150 units of gas were used for a total bill of $76. What is the unit cost of both types of energy?
Thank you so much !

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

During one month, a homeowner used 500 units of electricity and 100 units of gas for a total utility bill of $88. The next month, 400 units of electricity and 150 units of gas were used for a total bill of $76. What is the unit cost of both types of energy?
---
Equations:
500e + 100g = 88
400e + 150g = 76
----
Multiply thru the 1st equation by (3/2) to get
750e + 150g = 132
400e + 150g = 76
------
Subtract and solve for "e":
350e = 56
e = $0.16 or 16 cents per unit for electricity
----
Solve for "g":
500e + 100g = 88
500*0.16 + 100g = 88
80 + 100g = 88
100g = 8
g = $0.08 or 8 cents per unit for gas
===========================================
Cheers,
Stan H.
=========

Answer by mananth(16949)

let unit cost of electricity be $x
let unit cost of gas be $y
500x+100y=88.............1
400x+150y=76.............2

multiply (1)by -3
Multiply (2) by 2
-1500x-300y=-264
800x+300y=152

Add the two equations
-700x=-112
/-700
x=0.16 electricity unit cost 16 cents

plug value of x in (1)

500x+100y=88
80+100y =88
100y=88-80
100y=8
y=0.08 gas unit cost 8 cents